OpenStudy (anonymous):

What is the solution to the rational equation x over 2x minus 1 plus 1 over 4 equals 2 over 2x minus 1 ?

jhonyy9 (jhonyy9):

x 1 2 ----- + ---- = ------- so the common denominator will be 4(2x-1) 2x-1 4 2x -1 4x +2x-1 8 ----------- = --------- 4(2x-1) 4(2x-1) 6x-1 8 ---------- = -------- make the restriction for x so 2x-1 =0 ,result x=1/2 4(2x-1) 4(2x-1) so what mean that x not can being equal 1/2 6x-1=8 6x = 9 x= 9/6 x = 3/2